# Radio-loudness statistics of quasars from Quaia–VLASS

### v1

@@ Abstract
Context. Quasars are objects of high interest for applications in extragalactic astrophysics, cosmology, and astrometry. One of their useful qualities is owed to their relativistic jets: they can be radio-loud. However, the fraction of radio-loud vs. radio-quiet quasars is subject to ongoing investigations, where the statistical power is limited by the low number of known quasars with radio counterparts.
Aims. In this analysis, we revisited the radio-loudness statistics of quasars by significantly expanding the pool of known sources. Our main goal was to create a new, value-added catalogue of quasars with information about their extinction-corrected magnitudes, radio flux density, possible contamination levels, and other data flags, besides their sky coordinates and photometric redshifts.
Methods. We cross-matched the optical Quaia catalogue of about $1.3$ million quasars (selected from the Gaia data set) with $1.9$ million sources from the Very Large Array Sky Survey (VLASS) radio catalogue. We explored different thresholds for the matching radius, balancing the completeness and purity of the resulting Quaia–VLASS catalogue, and found that $1.5\arcsec$ is a sufficient choice.
Results. Our main finding is that the radio-loud fraction of quasars is in good agreement with previous works ($<10\%$), and there is no significant large-scale pattern in radio-loudness across the sky. The exact estimate depends on the $G$-band magnitude limit, and we observed some weak trends with redshift and absolute optical magnitude, possibly indicating remnant systematic effects in our data sets. The cross-matched Quaia–VLASS catalogue with 43,650 sources is available to the public for future analyses.
Conclusions. This latest census of QSOs with radio counterparts will facilitate further investigations of the dichotomy of radio-loud and radio-quiet quasars, and it may also support other lines of statistical investigations using quasars in cosmology and astrophysics.
Key Words.:
catalogs – surveys – galaxies: active – quasars: general – radio continuum: galaxies – large-scale structure of Universe
@@ 1 Introduction
Active galactic nuclei (AGN) are among the most powerful and dynamic objects in the Universe (e.g. Padovani et al., 2017), usually outshining their host galaxies. They can be found over a wide range of distances from the local Universe up to extremely high redshifts ($z\gtrsim 10$, Goulding et al., 2023). AGN play an important role in astrophysics, astrometry, and cosmology. They are located at the centres of galaxies and powered by accretion onto supermassive black holes (SMBHs) which range from millions to billions of solar masses. As material spirals into the SMBH, it heats up and emits huge amounts of energy across the entire electromagnetic spectrum, from radio to $\gamma$-rays.
Members of a small but spectacular subset of AGN exhibit intense radio emission originating from synchrotron radiation of relativistic charged particles moving in magnetized plasma jets launched from the close vicinity of SMBHs (e.g. Urry & Padovani, 1995; Blandford et al., 2019). Studying these so-called radio-loud, or rather jetted (Padovani, 2017) AGN is crucial for several reasons. Their jets are among the most extreme particle accelerators in the Universe, giving unique insight into relativistic effects, strong magnetic fields, and high-energy particle interactions. The radiative and mechanical feedback of AGN jets can influence the evolution of their host galaxies by regulating star formation (e.g. Fabian, 2012; Hardcastle & Croston, 2020).
Figure 1: The distribution of the $1.1$ million Quaia sources that fall within the joint coverage with the VLASS survey ($\delta>-40\degr$). The dust obscuration in the Milky Way disc results in the detection of fewer QSOs in the optical.
@@
A subclass of AGN, quasars (originally meaning quasi-stellar radio sources) were discovered as distant extragalactic objects in the early 1960’s (Schmidt, 1963). Not much later, it turned out that the majority of the so-called quasi-stellar objects (QSOs, also called quasars for simplicity) are not strong radio emitters (Sandage, 1965). Nowadays, the number of known quasars with measured spectroscopic redshifts exceeds 1 million (Flesch, 2023). The fraction of radio-loud quasars among the whole population is known to be $\lesssim 10\%$ (e.g. Ivezić et al., 2002; Kellermann et al., 2016).
@@
There are various working definitions of radio loudness in the literature. One can use the radio-to-optical flux density ratios as originally proposed by Schmidt (1970). Perhaps the most widely used definition is based on the ratio of the rest-frame $5$-GHz radio and 4400-Å optical flux densities (Kellermann et al., 1989). If this ratio $R$ exceeds 10 then the object is considered radio-loud. Sramek & Weedman (1980) used 2500-Å optical flux densities. Given the well-known flux density variability of quasars and the fact that optical and radio measurements are rarely contemporaneous, $R$ values themselves can be uncertain.
@@
The above two definitions are broadly consistent with each other for the quasar class as a whole (see a recent example in Krezinger et al., 2024). Alternatively, dictated by practical aspects such as the availability of measurements, it is customary to use other radio and optical bands, e.g. $1.4$ GHz in the radio and the $i$-band in the optical (like in Ivezić et al., 2002; Baloković et al., 2012). Note that Terashima & Wilson (2003) defined radio loudness as the ratio of radio ($5$ GHz) and X-ray ($2-10$ keV band) luminosities that can be applied for heavily obscured AGN. A fundamentally different approach is defining radio loudness by setting a threshold based on the (monochromatic) radio luminosity (e.g. Peacock et al., 1986; Miller et al., 1990). Interestingly, the two types of definitions provide generally similar classifications, as discussed in Appendix C of Ivezić et al. (2002).
@@
Figure 2: The distribution of the $1.9$ million VLASS2 QL sources after applying quality flags (see Table 1). Unlike for optical QSOs, the radio coverage is virtually unaffected by the Galactic disc.
@@
The ‘classical’ definition of radio loudness based on radio and optical flux density ratios is the best applicable for quasars where the optical emission is dominantly related to the accretion disk while the radio emission comes mainly from the jets (Padovani, 2017). Most early publications found a bimodal distribution of radio-loud and radio-quiet (in fact, not totally free from radio emission, rather radio-weak) quasars (e.g. Strittmatter et al., 1980; Kellermann et al., 1989; Miller et al., 1990; Ivezić et al., 2002). While the debate about this dichotomy is not settled yet, a consensus seems to emerge that radio loudness has a continuous distribution with no sharp gap between radio-loud and radio-quiet sources (see a review in Kellermann et al., 2016).
@@
For a long time, studies of radio loudness were naturally limited to small samples of relatively bright AGN. The emergence of sensitive large radio and optical sky surveys (e.g. Ivezić et al., 2002; Gürkan et al., 2019; Arnaudova et al., 2024) made substantially larger samples available for investigations. Here we revisit the problem with modern optical and radio surveys, and perform various statistical tests of radio-loudness of quasars.
@@
The paper is organized as follows. In Sect. 2, we introduce our observational data sets and the analysis methodology. Then, Sect. 3 contains our observational results, followed by a summary of our main conclusions in Sect. 4.
@@ 2 Data sets and methodology
The main goal of this work is to cross-match a large sample of optically detected quasars with a catalogue of radio sources, mainly following the footsteps of Ivezić et al. (2002), concentrating on the radio-loudness statistics. In this section, we present the input data sets and the technical details of the cross-matching process.
@@
For the optical quasar catalogue, we chose the recently compiled Quaia (Storey-Fisher et al., 2024) because of its large sky coverage, high completeness, and high astrometric accuracy. For the radio data, we relied on the Very Large Array Sky Survey (VLASS, Lacy et al., 2020) which is superior to all previous large sky surveys in terms of positional accuracy, angular resolution, and sensitivity.
2.1 Quasar catalogue
@@
For our cross-matched sample, we used the Quaia quasar catalogue (Storey-Fisher et al., 2024), culled from the Gaia space telescope’s data set (Gaia Collaboration et al., 2016). The QSOs of the Quaia catalogue are distinguished from the other types of sources by cross-matching the Gaia Data Release 3 (DR3) QSO candidates (Gaia Collaboration et al., 2023b, a; Delchambre et al., 2023) with infrared sources from the Wide-field Infrared Survey Explorer (WISE, Lang, 2014), applying colour cuts within the GaiaWISE photometry parameter space and cutting out sources with high proper motions.
@@
More precisely, we used the Quaia catalogue restricted in magnitude to $G<20.5$, containing about $1.3$ million QSOs, which yields high completeness of quasars within the Gaia sky coverage. There is also a conservative Quaia catalogue, a subsample of the larger one, with a constraint of $G<20.0$, with the benefit of the latter constraint yielding a purer QSO sample with a higher fidelity of the sources being QSOs. We decided to use the more complete catalogue with $G<20.5$, because its contaminants are expected to be mostly stars and should have negligible emission in the radio.
Another Quaia data product that we used is the so-called selection function, which indicates the QSO completeness due to various systematic effects (see Storey-Fisher et al., 2024, for details). The sky coverage of the Quaia catalogue that we consider in this analysis is displayed in Fig. 1.
@@
Figure 3: Left: the number of cross-matched sources as a function of a search radius parameter in arcseconds. We performed a robustness test using randomly perturbed sky coordinates, and found that a search radius of $1.5\arcsec$ results in a low level of contamination ($<0.1\%$) from false matches (right). We thus choose this threshold for our cross-matched Quaia–VLASS sample.
@@ 2.2 Radio sources
The other data set that we used is based on the ongoing VLASS (Lacy et al., 2020) at 3 GHz (S-band), conducted with the Very Large Array in New Mexico, in particular the VLASS Epoch 2 Quick Look (VLASS2 QL) catalogue. Unlike Quaia, the VLASS2 QL catalogue does not cover the full sky, and its 3.0 million radio sources are distributed at declinations $\delta>-40\degr$.
Further, VLASS is conducted in a single band, making it more difficult to deduct the class of the detected objects than for optical surveys with multiple bands, e.g. by segmentation in colour–colour space. Hence, cross-matching VLASS with another large-scale catalogue like Quaia is a straightforward choice for sampling radio-loud extragalactic objects. The sky area density of the VLASS2 QL catalogue, with applied quality flags, is displayed in Fig. 2.
@@
For our purposes, we applied the VLASS data flags explicitly recommended by the user manual (Gordon et al., 2023) to restrict ourselves to the most reliable data. We used the VLASS catalogue flag setting Duplicate_flag¡2 to get rid of any duplicated radio sources, originating from overlaps of the sky patches observed in the survey.
@@
Source catalogues $N_{\mathrm{sources}}$ Radio-loud
Quaia, full ($G<20.5$) 1,295,502 -
Quaia, $\delta>-40^{\circ}$ cut 1,084,545 -
VLASS2 QL 2,995,025 -
VLASS2 QL, flags applied 1,873,524 -
Quaia–VLASS (all) 43,650 $3.83\%\pm 0.02\%$
$18.514.0<G<$ 18.5 cut 9,395 $8.04\%\pm 0.09\%$
$19.518.5\leq G<$ 19.5 cut 15,763 $4.71\%\pm 0.04\%$
$20.519.5\leq G<$ 20.5 cut 18,492 $2.80\%\pm 0.02\%$
 [^1] Top: the number of sources contained in the data sets that we used. Bottom: The radio-loud fraction of sources in the full Quaia–VLASS cross-match and in subsets due to different cuts in $G$ optical magnitude (errors are based on Poisson statistics).
@@ 2.3 Cross-matching
Figure 5: The extent of our Quaia–VLASS sample along the $G$ (optical) and $t$ (radio) bands. Radio-loud and radio-quiet sources are separated by the black line. We defined three bins in the $G$ band, which we later analyse separately in the rest of the paper when looking for brightness-dependent trends in radio loudness.
@@
In Fig. 3, we present our findings in the context of the fraction of possibly false matches. In the left panel, $N_{\mathrm{cross}}$ is the cumulative number of cross-matches between Quaia and VLASS as a function of the coordinate separation up to the corresponding $\Delta\theta$ value. We also display $N_{\rm false}$, calculated with a method used previously by Orosz & Frey (2013). $N_{\rm false}$ measures a cross-matched fraction between VLASS and Quaia*, where Quaia* is a surrogate of the Quaia QSO catalogue with all sources rotated randomly by the same angle in right ascension. This ensures that $N_{\rm false}$ from cross-matching Quaia* and VLASS only contains unphysical chance alignments. Moreover, $N_{\rm false}$ should approximate the chance alignments between Quaia and VLASS very well, because Quaia* has the same source density as Quaia. The curve of $N_{\mathrm{corrected}}=N_{\mathrm{cross}}-N_{\rm false}$ demonstrates a saturation in the number of physical cross-matches at $\Delta\theta\approx 1.5\arcsec$ (displayed as a grey band).
Using the above estimate of the contamination, in Fig. 3 (right panel) we show the fraction of chance alignments as a function of $\Delta\theta$, which can be expressed as
@@
Figure 4 shows a histogram of the Quaia–VLASS angular separations up to $\Delta\theta=1.5\arcsec$, indicating that most cross-matches are identified with $\Delta\theta<0.5\arcsec$ separation, with a peak at about $\Delta\theta\approx 0.07\arcsec$ (see inset of Fig. 4). We thus conclude that the construction of our cross-matched sample guarantees high completeness.

### v2

@@ Abstract
Quasars are objects of high interest in extragalactic astrophysics, cosmology, and astrometry. One of their useful qualities is their potential radio loudness. However, the fraction of radio-loud vs. radio-quiet quasars is subject to ongoing investigations, where the statistical power is limited by the low number of known quasars with radio counterparts. In this analysis, we revisited the radio loudness statistics of quasars by significantly expanding the pool of known sources. Our main goal was to create a new, value-added quasar catalog with information about their extinction-corrected magnitudes, radio flux density, possible contamination levels, and other flags, besides their sky coordinates and photometric redshifts. We cross-matched the optical Quaia catalog of about $1.3$ million quasars (from the Gaia data) with $1.9$ million sources from the Very Large Array Sky Survey (VLASS) radio catalog. We explored different thresholds for the matching radius, balancing completeness and purity of the resulting Quaia–VLASS catalog, and found $1.5\arcsec$ a sufficient choice. Our main finding is that the quasar radio-loud fraction is in good agreement with previous works ($10\%$), and there is no significant large-scale sky pattern in radio loudness. The exact estimate depends on the $G$-band magnitude limit, and we observed weak trends with redshift and absolute optical magnitude, possibly indicating remnant systematic effects in our data. The cross-matched Quaia–VLASS catalog with 43,650 sources and accompanying code are available at 10.5281/zenodo.16035690. This latest census of QSOs with radio counterparts will facilitate further investigations of the dichotomy of radio-loud and radio-quiet quasars, and it may also support other lines of research.
Catalogs (205) — Surveys (1671) — Active galaxies (17) — Quasars (1319) — Radio loud quasars (1349) — Large-scale structure of the universe (902)
@@ 1 Introduction
Active galactic nuclei (AGN) are among the most powerful and dynamic objects in the Universe (e.g. Padovani et al., 2017; Schindler et al., 2019; Jorstad et al., 2022), some of which can outshine their host galaxies. They can be found over a wide range of distances from the local Universe up to extremely high redshifts ($z\gtrsim 10$, Goulding et al., 2023). AGN play an important role in astrophysics, astrometry, and cosmology. They are located at the centres of galaxies and powered by accretion onto supermassive black holes (SMBHs) which range from millions to billions of solar masses. As material spirals into the SMBH, it heats up and emits huge amounts of energy across the entire electromagnetic spectrum, from radio to $\gamma$-rays (e.g. Padovani, 2017a).
Members of a small but spectacular subset of AGN exhibit intense radio emission originating from synchrotron radiation of relativistic charged particles moving in magnetized plasma jets launched from the close vicinity of SMBHs (e.g. Urry & Padovani, 1995; Blandford et al., 2019). Studying these so-called radio-loud, or rather jetted (Padovani, 2017b) AGN is crucial for several reasons. Their jets are among the most extreme particle accelerators in the Universe, giving unique insight into relativistic effects, strong magnetic fields, and high-energy particle interactions. The radiative and mechanical feedback of AGN jets can influence the evolution of their host galaxies by regulating star formation (e.g. Fabian, 2012; Hardcastle & Croston, 2020).
@@
A subclass of AGN, quasars (originally meaning quasi-stellar radio sources) were discovered as distant extragalactic objects in the early 1960s (Schmidt, 1963). Not much later, it turned out that the majority of the so-called quasi-stellar objects (QSOs) are not strong radio emitters (Sandage, 1965). Nowadays, the number of known quasars with measured spectroscopic redshifts exceeds 1 million (Flesch, 2023). The fraction of radio-loud quasars among the whole population is known to be $\lesssim 10\%$ (e.g. Ivezić et al., 2002; Kellermann et al., 2016).
@@
There are various working definitions of radio loudness in the literature (Kellermann et al., 1989; Ivezić et al., 2002; Baloković et al., 2012; Macfarlane et al., 2021; Yue et al., 2024). One can use the radio-to-optical flux density ratios as originally proposed by Schmidt (1970). Perhaps the most widely used definition is based on the ratio of the rest-frame $5$-GHz radio and 4400-Å optical flux densities (Kellermann et al., 1989). If this ratio $R$ exceeds 10 the object is considered radio-loud. Sramek & Weedman (1980) used 2500-Å optical flux densities. Given the well-known flux density variability of quasars and the fact that optical and radio measurements are rarely contemporaneous, $R$ values themselves can be uncertain.
@@
The above two definitions are broadly consistent with each other for the quasar class as a whole (see a recent example in Krezinger et al., 2024). Alternatively, dictated by practical aspects such as the availability of measurements, it is customary to use other radio and optical bands, e.g. $1.4$ GHz in the radio and the $i$-band in the optical (like in Ivezić et al., 2002; Baloković et al., 2012). Note that Terashima & Wilson (2003) defined radio loudness as the ratio of radio ($5$ GHz) and X-ray (2102-10 keV band) luminosities that can be applied for heavily obscured AGN. Similarly, Klindt et al. (2019) introduced a radio-loudness metric based on the ratio of rest-frame $1.4$ GHz radio luminosity to $6~\mathrm{\mu m}$ mid-infrared luminosity, which is less affected by dust extinction than traditional optical-based definitions. A fundamentally different approach is defining radio loudness by setting a threshold based on the (monochromatic) radio luminosity (e.g. Peacock et al., 1986; Miller et al., 1990). Interestingly, the later definition and the one by Terashima & Wilson (2003) provide generally similar classifications, as discussed in Appendix C of Ivezić et al. (2002).
@@
Figure 1: On the left the distribution of the $1.1$ million Quaia sources that fall within the joint coverage with the VLASS survey ($\delta>-40\arcdeg$). The dust obscuration in the Milky Way disc results in the detection of fewer QSOs in the optical. On the right the distribution of the $1.9$ million VLASS2 QL sources after applying quality flags (see Table 1). Unlike for optical QSOs, the radio coverage is virtually unaffected by the Galactic disc.
@@
The ‘classical’ definition of radio loudness based on radio and optical flux density ratios is the best applicable for quasars where the optical emission is dominantly related to the accretion disk while the radio emission comes mainly from the jets (Padovani, 2017b). Nonetheless, Macfarlane et al. (2021) show that for radio-quiet quasars the majority of the radio emission can come from star formation, and Panessa et al. (2022) show that other forms of emission can contribute significantly to the radio flux of radio-quiet quasars, e.g. AGN winds. Most early publications found a bimodal distribution of radio-loud and radio-quiet (in fact, not totally free from radio emission, rather radio-weak) quasars (e.g. Strittmatter et al., 1980; Kellermann et al., 1989; Miller et al., 1990; Ivezić et al., 2002). While the debate about this dichotomy is not settled yet, a consensus seems to emerge that radio loudness has a continuous distribution with no sharp gap between radio-loud and radio-quiet sources (see a review in Kellermann et al., 2016; Gürkan et al., 2019; Macfarlane et al., 2021; Arnaudova et al., 2024; Yue et al., 2024).
@@
For a long time, studies of radio loudness were naturally limited to small samples of relatively bright AGN. The emergence of sensitive large radio and optical sky surveys (e.g. Ivezić et al., 2002; Gürkan et al., 2019; Arnaudova et al., 2024) made substantially larger samples available for investigations (see Table 2). Here we revisit the radio loudness and fraction of radio-loud QSOs with two modern optical and radio surveys and we test statistically the dependence of radio loudness on redshift, absolute optical magnitude and selection function (later explained in Section 2.1).
@@
The paper is organized as follows. In Section 2, we introduce our observational data sets and the analysis methodology. Then, Section 3 contains our observational results, followed by a summary of our main conclusions in Section 4.
[…]
@@ 2 Data Sets and Methodology
The main goal of this work is to cross-match a large sample of optically detected quasars with a catalog of radio sources, mainly following the footsteps of Ivezić et al. (2002), concentrating on the radio loudness statistics. In this Section, we present the input data sets and the technical details of the cross-matching process.
@@
For the optical quasar catalog, we chose the recently compiled Quaia (Storey-Fisher et al., 2024) because of its large sky coverage, high completeness, and high astrometric accuracy. There are several recent radio surveys available for the given goals of this work, e.g. the Very Large Array Sky Survey (VLASS, Lacy et al., 2020) and the LOFAR Two-metre Sky Survey (LoTSS) (Shimwell et al., 2017). The latter operates at 144 MHz and offers high sensitivity for synchrotron-dominated emission, making it particularly valuable for studying extended radio AGN and jet structures. We decided to use VLASS, which operates at a higher frequency (3 GHz) and is superior to LoTSS in positional accuracy, angular resolution and size of covered sky area.
2.1 Quasar Catalog
@@
For our cross-matched sample, we used the Quaia quasar catalog (Storey-Fisher et al., 2024), culled from the Gaia space telescope’s data set (Gaia Collaboration et al., 2016). The QSOs of the Quaia catalog are distinguished from the other types of sources by cross-matching the Gaia Data Release 3 (DR3) QSO candidates (Gaia Collaboration et al., 2023a, b; Delchambre et al., 2023) with infrared sources from the Wide-field Infrared Survey Explorer (WISE, Lang, 2014), applying color cuts within the Gaia–WISE photometry parameter space and cutting out sources with high proper motions. The Quaia redshifts, that we use through out this work, are estimated by Storey-Fisher et al. (2024) with a k-Nearest-Neighbors (kNN) model, using as training features colors from the Gaia and WISE photometry, the Gaia $G$-band magnitude, the dust reddening E(BV)E(B-V) at the location of the sources, and cruder estimates for the redshifts of the QSOs from Gaia Collaboration et al. (2023b). The labels for the kNN training are SDSS redshift values, for the subset of QSOs with available measurements.
@@
More precisely, we used the Quaia catalog restricted in magnitude to $G<20.5$, containing about $1.3$ million QSOs, which yields high completeness of quasars within the Gaia sky coverage. There is also a conservative Quaia catalog, a subsample of the larger one, with a constraint of $G<20.0$, with the benefit of the latter constraint yielding a purer QSO sample with a higher fidelity of the sources being QSOs. We decided to use the more complete catalog with $G<20.5$, because its contaminants are expected to be mostly stars and should have negligible emission in the radio.
Another Quaia data product that we used is the so-called selection function. It is a healpix (Gorski et al., 2005) all-sky map with pixel values in the range $[0,1]$. The values indicate to what extent the QSO count completeness in the particular pixel has been diminished by the most severe systematics – dust, the source density of the parent surveys, and the scan patterns of the parent surveys. Hence the selection function is a useful tool to correct for these systematics in a statistical analysis. The sky coverage of the Quaia catalog that we consider in this analysis is displayed in Figure 1.
The Quaia catalog has several limitations, which have to be taken into account. On the one side, the redshift estimation for Quaia is spectrophotometric and superior to the initial Gaia redshift estimates, but inferior to spectroscopic ones, like those from SDSS (Lyke et al., 2020). Users should consider the redshift uncertainties and the presence of catastrophic outliers in Quaia. The Quaia selection function carries limitations, too: diminished accuracy close to the Galactic disc, the inevitability to mask out Galactic regions for some tasks, dependence on right ascension and declination, but not redshift, assumption of isotropy, which is contested (Stewart & Sciama, 1967; Secrest et al., 2021). For a deeper discussion of the Quaia limitations, see Storey-Fisher et al. (2024). Our work is independent of some of the limitations, mostly regarding cosmological applications. For instance, the lower number density of Quaia compared to other surveys, like SDSS, is not a concern in this work, nor are isotropy assumptions or the lack of homogeneity in the distribution of Quaia or VLASS sources. The radio-loud fractions (RLFs) presented in Table 1 are not influenced by limitations in Quaia’s redshift or selection function values, nor is the cross-matching of Quaia and VLASS, since it depends only on the right ascension and declination values. Nevertheless, potential Quaia miss-classifications of galaxies as QSOs would lead to impurities in our data product of Quaia–VLASS quasar cross-matches. Evidence for such impurities is discussed in Section 3.5. A further limitation of Quaia is that the uncertainty in the redshifts makes the search for correlation of properties with redshift and $G_{\mathrm{abs}}$ more difficult, as discussed in Section 3.5. Not mentioned by Storey-Fisher et al. (2024), the $G$-band is not extinction-corrected, which we take care of in Section 2.4 and observe a strong improvement of the extinction bias.
@@
Figure 2: Left: the number of cross-matched sources as a function of a search radius parameter in arcseconds. We performed a robustness test using randomly perturbed sky coordinates, and found that a search radius of $1.5\arcsec$ results in a low level of contamination ($0.1\%$) from false matches (right). We thus choose this threshold for our cross-matched Quaia–VLASS sample.
@@ 2.2 Radio Sources
The other data set that we used is based on the ongoing VLASS (Lacy et al., 2020) at 3 GHz (S-band), conducted with the Very Large Array in New Mexico, in particular the VLASS Epoch 2 Quick Look (VLASS2 QL) catalog. Unlike Quaia, the VLASS2 QL catalog does not cover the full sky, and its 3.0 million radio sources are distributed at declinations $\delta>-40\arcdeg$.
Further, VLASS is conducted in a single band, making it more difficult to deduce the class of the detected objects than for optical surveys with multiple bands, e.g. by segmentation in color–color space. Hence, cross-matching VLASS with another large-scale catalog like Quaia is a straightforward choice for sampling radio-loud extragalactic objects. The sky area density of the VLASS2 QL catalog, with applied quality flags, is displayed on the right in Figure 1.
@@
Source catalogs $N_{\mathrm{sources}}$ $N_{\mathrm{loud}}$ $N_{\mathrm{Quaia}}$ RLF ($N_{\mathrm{loud}}/N_{\mathrm{Quaia}}$)
[…]
     14.0 <G<<G< 18.5 9395 7431 92397 8.04% $\pm$ 0.09%
     18.5 $\leq G<$ 19.5 15763 15576 330779 4.71% $\pm$ 0.04%
     19.5 $\leq G<$ 20.5 18492 18489 661369 2.8% $\pm$ 0.02%
@@ 2.3 Cross-matching
Figure 4: The extent of our Quaia–VLASS sample along the $G$ (optical) and $t$ (radio) bands. Radio-loud and radio-quiet sources are separated by the black line. We defined three bins in the $G$ band, which we later analyse separately in the rest of the paper when looking for brightness-dependent trends in radio loudness. The dashed line shows the $t=16.6$ sensitivity limit, as inferred from the RMS noise limit stated in Lacy et al. (2022), which agrees reasonably well with the lack of sources in the figure at $t>16.6$. The Quaia catalog has a limit at $G=20.5$ as imposed by Storey-Fisher et al. (2024). The region within which there is sensitivity in $G$ and sources would be radio-loud, but lies beyond the sensitivity limit in $t$, is labelled “Indeterminate region”.
@@
In Figure 2, we present our findings in the context of the fraction of possibly false matches. In the left panel, $N_{\mathrm{cross}}$ is the cumulative number of cross-matches between Quaia and VLASS as a function of the coordinate separation up to the corresponding $\Delta\theta$ value. We also display $N_{\rm false}$, calculated with a method used previously by Orosz & Frey (2013). $N_{\rm false}$ measures a cross-matched fraction between VLASS and Quaia*, where Quaia* is a surrogate of the Quaia QSO catalog with all sources rotated randomly by the same angle in right ascension. This ensures that $N_{\rm false}$ from cross-matching Quaia* and VLASS only contains unphysical chance alignments. Moreover, $N_{\rm false}$ should approximate the chance alignments between Quaia and VLASS very well, because Quaia* has the same source density as Quaia. The curve of $N_{\mathrm{corrected}}=N_{\mathrm{cross}}-N_{\rm false}$ demonstrates a saturation in the number of physical cross-matches at $\Delta\theta\approx 1.5\arcsec$ (displayed as a grey band).
Using the above estimate of the contamination, in Figure 2 (right panel) we show the fraction of chance alignments as a function of $\Delta\theta$, which can be expressed as
@@
Figure 3 shows a histogram of the Quaia–VLASS angular separations up to $\Delta\theta=1.5\arcsec$, indicating that most cross-matches are identified with $\Delta\theta<0.5\arcsec$ separation, with a peak at about $\Delta\theta\approx 0.07\arcsec$ (see inset of Figure 3). We thus conclude that the construction of our cross-matched sample guarantees high completeness. The number of the Quaia–VLASS cross-matches in different $G$ bins and in total can be looked up in Table 1.